# Intrinsic Wrapped Gaussian Process Regression Modeling for Manifold-valued Response Variable

### v1

@@ 2.3 Intrinsic wrapped Gaussian process regression
${\scriptsize{\binom{y_{t^{*}}}{\boldsymbol{y}}\sim\mathcal{N}_{\mathcal{M}}\left(\binom{\mu_{t^{*}}}{\boldsymbol{\mu}},\left(\begin{array}[]{cc}\mathbf{E}^{\top}\boldsymbol{K}_{\mathbf{E},t^{*}}^{**}\mathbf{E},&\mathbf{E}^{\top}\boldsymbol{K}_{\mathbf{E},t^{*}}^{*}(\mathbf{E}\otimes\boldsymbol{\boldsymbol{I}}_{n\times n})\\ (\mathbf{E}^{\top}\boldsymbol{K}_{\mathbf{E},t^{*}}^{*}(\mathbf{E}\otimes\boldsymbol{\boldsymbol{I}}_{n\times n}))^{\top},&(\mathbf{E}\otimes\boldsymbol{\boldsymbol{I}}_{n\times n})^{\top}(\boldsymbol{K}_{\mathbf{E},t^{*}}+\sigma^{2}\boldsymbol{\boldsymbol{I}}_{nD\times nD})(\mathbf{E}\otimes\boldsymbol{\boldsymbol{I}}_{n\times n})\end{array}\right)\right),}}$ (17)
where $\mathcal{N}_{\mathcal{M}}$ is the wrapped gaussian distribution. The remaining procedure on computing estimation and prediction is the same as the one in Subsection 2.3, where $(\mathbf{E}\otimes\boldsymbol{\boldsymbol{I}}_{n\times n})^{\top}\boldsymbol{K}_{\mathbf{E},t^{*}}(\mathbf{E}\otimes\boldsymbol{\boldsymbol{I}}_{n\times n})$ is replaced with $(\mathbf{E}\otimes\boldsymbol{\boldsymbol{I}}_{n\times n})^{\top}(\boldsymbol{K}_{\mathbf{E},t^{*}}+\sigma^{2}\boldsymbol{\boldsymbol{I}}_{nD\times nD})(\mathbf{E}\otimes\boldsymbol{\boldsymbol{I}}_{n\times n})$. Theorem 2.8 below proves information consistency of the prediction obtained by the model (14). Suppose the underlying true vector field $\boldsymbol{\gamma}(t)$ along the BPF $\mu(t)$ can be represented by real-valued function vectors $\boldsymbol{\gamma}_{\mathbf{E}}$ based on orthogonal coordinate frame $\{\mathbf{E}_{t}\}_{t\in\mathbf{T}}$, where $\Gamma_{t\rightarrow s}(\mathbf{E}_{t})=\mathbf{E}_{s}$ for any $t,s\in\mathbf{T}$. Without loss of generality, we consider that a special setting of $k_{\mathbf{E},t^{*}}$, i.e, the coordinates are independent of each other, i.e, $k_{\mathbf{E},t^{*}}\left(\cdot,\cdot\right)=\text{diag}(k_{\mathbb{R}^{Q}}\left(\cdot,\cdot,\boldsymbol{\theta}_{1}\right),\cdots,k_{\mathbb{R}^{Q}}\left(\cdot,\cdot,\boldsymbol{\theta}_{D}\right))$. Then, assume that the random element $\{\boldsymbol{\gamma}_{\mathbf{E},d}\},d=1,2,\cdots,D$ has a Gaussian process prior with mean zero and bounded covariance function $k_{\mathbb{R}^{Q}}\left(\cdot,\cdot,\boldsymbol{\theta}_{d}\right)$ for any covariate $\boldsymbol{x}_{t}$. Suppose the $k_{\mathbb{R}^{Q}}\left(\cdot,\cdot,\boldsymbol{\theta}_{d}\right)$ is continuous in $\boldsymbol{\theta}_{d}$, further assume the estimator $\widehat{\boldsymbol{\theta}}_{d}$ converges to $\boldsymbol{\theta}_{d}$ as $n\rightarrow\infty$ almost surely, and the BPF $\mu(t)$ is known.
@@
Under the above assumptions, let the reproducing kernel Hilbert space norm is bounded and the expected regret term $E\left(\log\left|\boldsymbol{I}_{nD\times nD}+\sigma^{-2}\boldsymbol{K}_{\mathbf{E}}\right|\right)=o(n)$, where $\boldsymbol{K}_{\mathbf{E}}=\left(k_{\mathbf{E},t^{*}}\left(\boldsymbol{x}_{t_{i}},\boldsymbol{x}_{t_{j}}\right)\right)_{1\leq i,j\leq n}$ is the representation of the covariance matrix over $\{\boldsymbol{x}_{t_{i}}\}_{i=1}^{n}$ under the orthogonal coordinate frames $\{\mathbf{E}_{t}\}_{t\in\mathbf{T}}$, and $\sigma^{2}$ is the variance of the measurement error. Then, the estimator $\widehat{\boldsymbol{\gamma}}_{\mathbf{E},d}$ is information consistent to the true ${\boldsymbol{\gamma}}_{\mathbf{E},d}$, which means an estimator that guarantees convergence to the true parameter value as the sample size grows.
The proof is given in Appendix F. Furthermore, we also consider the consistency of the posterior distribution. This is shown in Theorem 2.9 and the proof is given in Appendix G.
@@
Suppose the underlying true vector field $\boldsymbol{\gamma}(t)$ along the BPF $\mu(t)$ can be represented by real-valued function vectors $\boldsymbol{\gamma}_{\mathbf{E}}$ based on orthogonal coordinate frame $\{\mathbf{E}_{t}\}_{t\in\mathbf{T}}$, where $\Gamma_{t\rightarrow s}(\mathbf{E}_{t})=\mathbf{E}_{s}$ for any $t,s\in\mathbf{T}$. Without loss of generality, we consider that the coordinates are independent of each other and the BPF $\mu(t)$ is known. Let ${\boldsymbol{\gamma}}_{\mathbf{E},d}:\mathbb{R}^{Q}\rightarrow\mathbb{R}$ be a function and $\widehat{v}_{n,d}(\boldsymbol{x}_{t})$ denotes the variance of the posterior distribution of the estimator $\widehat{\boldsymbol{\gamma}}_{E,d}$. Then, for every $d=1,2,\cdots,D$,
$\sup_{\boldsymbol{x}\in\mathbb{R}^{Q}}\widehat{v}_{n,d}(\boldsymbol{x}_{t})\longrightarrow 0,\quad n\rightarrow\infty,$
almost surely monotonically, as well as in $L^{p}$ for every $p\in[1,\infty)$. Furthermore, if ${\boldsymbol{\gamma}}_{\mathbf{E},d}$ lies in the RKHS of $k_{\mathbb{R}^{Q}}\left(\cdot,\cdot,\boldsymbol{\theta}_{d}\right)$, for all $\boldsymbol{x}_{t}\in\mathbb{R}^{Q}$ and $d=1,2,\cdots,D$,
$\widehat{\boldsymbol{\gamma}}_{\mathbf{E},d}(\boldsymbol{x}_{t})\xrightarrow{L^{2}}{\boldsymbol{\gamma}}_{\mathbf{E},d}(\boldsymbol{x}_{t}),\quad n\rightarrow\infty.$
@@
In Theorem 2.9, it needs the condition that $\mu(t)$ is known. But in practice, $\mu(t)$ is usually unknown and is estimated by using data. In this case, we have the following conclusion. Consider the same conditions as in Theorem 2.9 and suppose $\sup_{t\in\mathcal{T}}d_{\mathcal{M}}^{2}(\widehat{\mu}(t),\mu(t))=O_{P}\left(n^{-1}\right)$ holds. Then, for all $\boldsymbol{x}_{t}\in\mathbb{R}^{Q}$, we have $(\Gamma_{\widehat{\mu}(t)\rightarrow\mu(t)}(\widehat{\boldsymbol{\gamma}}(\boldsymbol{x}_{t})))_{\mathbf{E},d}\xrightarrow{L^{2}}{\boldsymbol{\gamma}}_{\mathbf{E},d}(\boldsymbol{x}_{t}),\quad n\rightarrow\infty$. Furthermore, from Theorem 6 of Lin and Yao [9], assuming that the basepoint function is an intrinsic mean curve, then the condition $\sup_{t\in\mathcal{T}}d_{\mathcal{M}}^{2}(\widehat{\mu}(t),\mu(t))=O_{P}\left(n^{-1}\right)$ holds.
@@ 3 Simulation studies
In this section, we conduct several numerical experiments to illustrate the performance of the proposed iWGPR model. We consider sample sizes of $N=30,50,100,150$ and two setups. The first one is a random sampling scheme (Random), where 80% of the $N$ observation points are randomly selected as training data and the remainings are the test data. The second setup is a sorted sampling scheme (Sort), in which the last five samples of the $N$ sample points are designated as the test data and the remainings are the training data. All simulation results are based on 100 independent repetitions, and the average and standard errors of the root mean square geodesic errors are computed. To evaluate the performance of the proposed method, we compare it with wrapped Gaussian process regression (WGPR) [23] and vector-valued Gaussian process regression (MGPR) [38].
Table 1: The average and standard errors of the root mean square geodesic errors of WGPR, iWGPR and MGPR on the SPD manifold $\mathcal{S}_{++}^{3}$ with S1.
Scheme $\boldsymbol{\theta}$ $n$ iWGPR WGPR MGPR
@@
Table 2: The average and standard errors of the root mean square geodesic errors of WGPR, iWGPR and MGPR on the SPD manifold $\mathcal{S}_{++}^{3}$ with S2.
Scheme $\boldsymbol{\sigma^{2}}$ $n$ iWGPR WGPR MGPR
@@
3.1 Symmetric positive definite matrices
Figure 3: Fitted covariance function via WGPR and iWGPR on the tangent space of manifold ${\mathcal{S}_{++}^{3}}$. (S1, $\theta_{1}=0.1$, $\theta_{2}=0.3$, $\theta_{3}=0.5$, $n=30$, sort scheme). The true covariance functions for the first, second, and third dimensions are shown on the left, arranged from top to bottom. The middle section presents the covariance function estimation results for each dimension using the iWGPR method, while the right side displays the estimation results using the WGPR method.
We first consider the symmetric positive definite matrices space $\mathcal{S}_{++}^{3}$ with affine-invariant metric. Let $\mu(t)$ be a geodesic curve
$\mu(t)=\operatorname{Exp}(\mu(0),\operatorname{Log}(\mu(0),(2.4360,-2.6465;-2.6465,8.4208)^{\top})t),$
where $\mu(0)=(1,0;0,1)$, the time domain is $t\in[0,1]$ and the predictor is time, i.e, $\boldsymbol{x}_{t}=t$. The initial frame $(1,0;0,0)$, $(0,1;1,0)$ and $(0,0;0,1)$ are utilized in the tangent space of $\mu(0)$. Let the coordinates be independent using the tangle space of $\mu(0)$ as initial tangle space to generate the covariance function. We consider the following two scenarios. (S1) We use the following Gaussian kernel function matrix as the intrinsic covariance function under the frame $\mathbf{E}$,
$k_{\mathbf{E},t^{*}}(\boldsymbol{x},\widetilde{\boldsymbol{x}}):=\begin{pmatrix}\exp({\frac{-\|\boldsymbol{x}-\widetilde{\boldsymbol{x}}\|_{2}^{2}}{\theta_{j1}}})&0&0\\ 0&\exp({\frac{-\|\boldsymbol{x}-\widetilde{\boldsymbol{x}}\|_{2}^{2}}{\theta_{j2}}})&0\\ 0&0&\exp({\frac{-\|\boldsymbol{x}-\widetilde{\boldsymbol{x}}\|_{2}^{2}}{\theta_{j3}}})\end{pmatrix},\quad\boldsymbol{x},\widetilde{\boldsymbol{x}}\in[0,1];\quad j=1,2,3,$
Average values and standard errorss of the root mean square geodesic errors from three methods, iWGPR, WGPR and MGPR, are shown in Tables 1-2, corresponding to scenarios S1 and S2 respectively. From both Tables, we see that iWGPR has the smallest root mean square geodesic errors, indicating that iWGPR has a significant improvement in predictive performance compared to WGPR and MGPR. In addition, MGPR has the largest root mean square geodesic errors, as it fails to account for the geometric information inherent in manifold-valued data during modeling, thus violating the intrinsic structure of the manifold. Although WGPR incorporates the local nonlinear structure of the manifold through exponential and logarithmic maps, it still disrupts the manifold’s intrinsic geometric structure by directly computing tangent vectors from different tangent spaces during modeling. As the sampling points become sparser, the proposed method still maintains excellent performance, whereas the performance of WGPR declines significantly.
We also investigate the ability of the proposed method to reconstruct the covariance function of wrapped Gaussian process. Figure 3 illustrates the estimated covariance function in the tangent space comared to the true covariance function with kernel parameters for covariance structure reflecting the mean values from 100 experiments. The heatmaps in Figures 3 indicate that the estimated covariance functions align more closely with the true one compared with WGPR.
3.2 Unit sphere
Table 3: The average and standard errors of the root mean square geodesic errors of WGPR, iWGPR and MGPR on the sphere manifold $\mathcal{S}^{2}$ with S3.
Scheme $\boldsymbol{\theta}$ $n$ iWGPR WGPR MGPR
Random (0.1, 0.5) 30 0.0261(0.1643) 0.0352(0.1736) 1.3527(0.5749)
50 0.0154(0.0939) 0.0217(0.1129) 1.2930(0.5639)
100 0.0128(0.1085) 0.0140(0.1089) 1.3730(0.4849)
150 0.0062(0.0967) 0.0062(0.0989) 1.0138(0.5182)
(0.08, 0.8) 30 0.0121(0.0918) 0.0242(0.1177) 1.8390(0.6729)
50 0.0101(0.0775) 0.0148(0.0857) 2.0183(0.5812)
100 0.0089(0.0829) 0.0140(0.0930) 1.4309(0.5819)
150 0.0075(0.0727) 0.0076(0.0877) 1.3302(0.4923)
Sort (0.1, 0.5) 30 0.1196(0.1478) 0.1657(0.1524) 1.6839(0.5389)
50 0.0709(0.1218) 0.0860(0.1273) 2.2283(0.4492)
100 0.0344(0.1033) 0.0476(0.1156) 2.1943(0.5784)
150 0.0145(0.0602) 0.0144(0.0604) 1.9920(0.6492)
(0.08, 0.8) 30 0.1334(0.2105) 0.1962(0.2132) 2.3819(0.4837)
50 0.0710(0.1184) 0.0716(0.2110) 2.2917(0.5181)
100 0.0431(0.1130) 0.0455(0.2069) 1.2839(0.5379)
150 0.0255(0.1821) 0.0255(0.1892) 1.2903(0.6839)
We now consider the unit sphere $\mathcal{S}^{2}$, where $\mathcal{S}^{2}$ is a compact nonlinear Riemannian submanifold of $\mathbb{R}^{3}$, i.e, the unit sphere $\mathcal{S}^{2}$ consists of points $(x,y,z)\in\mathbb{R}^{3}$ satisfying $x^{2}+y^{2}+z^{2}=1$. Let $\mu(t)$ be a geodesic curve $\operatorname{Exp}(\mu(0),\operatorname{Log}(\mu(0)(-1/\sqrt{3},1/\sqrt{3},-1/\sqrt{3})^{\top})t)$, where $\mu(0)=(0,0,1)^{\top}$ and the time domain is set to $t\in[0,1]$ and the predictor is time index, i.e, $\boldsymbol{x}_{t}=t$. The initial frame $(1,~{}0,~{}0)^{\top}$ and $(0,1,0)^{\top}$ are utilized in the tangent space of $\mu(0)$. Let the coordinates be independent and the tangle space of $\mu(0)$ be initial tangle space to generate the covariance function. We also consider two scenarios. (S3) Gaussian kernel function matrix as the intrinsic covariance function under the frame $\mathbf{E}$ is defined as,
$k_{\mathbf{E},t^{*}}(\boldsymbol{x},\widetilde{\boldsymbol{x}}):=\begin{pmatrix}\exp({\frac{-\left\|\boldsymbol{x}-\widetilde{\boldsymbol{x}}\right\|_{2}^{2}}{\theta_{j1}}})&0\\ 0&\exp({\frac{-\left\|\boldsymbol{x}-\widetilde{\boldsymbol{x}}\right\|_{2}^{2}}{\theta_{j2}}})\end{pmatrix},\quad\boldsymbol{x},\widetilde{\boldsymbol{x}}\in[0,1];\quad j=1,2,$
with parameters $(\theta_{11},\theta_{12})=(0.1,~{}0.5)^{\top}$ and $(\theta_{21},\theta_{22})=(0.08,~{}0.8)^{\top}$. (S4) $(\gamma_{1,\mathbf{E}}(t),\gamma_{1,\mathbf{E}}(t))$ is used as the vector-valued function under the frame $\mathbf{E}$, where $\gamma_{1,\mathbf{E}}(t)=1+0.05*t+\sin(t)/(t+0.001)+0.01*\varepsilon_{1,t}$ and $\gamma_{2,\mathbf{E}}(t)=1+\cos(t)+0.05*\varepsilon_{2,t}$. $\varepsilon_{1,t}$ and $\varepsilon_{2,t}$ are standard normal distributions.
Table 4: The average and standard errors of the root mean square geodesic errors of WGPR, iWGPR and MGPR on the sphere manifold $\mathcal{S}^{2}$ with S4.
Scheme $\boldsymbol{\sigma^{2}}$ $n$ iWGPR WGPR MGPR
Random (0.001,0.005) 30 0.0815(0.1045) 0.1059(0.1047) 1.4483(0.5389)
50 0.0728(0.0893) 0.1028(0.0927) 1.4767(0.6438)
100 0.0705(0.0952) 0.0762(0.0958) 1.5639(0.4920)
150 0.0270(0.0665) 0.0285(0.0693) 1.3830(0.5290)
Sort (0.001,0.005) 30 0.0304(0.0106) 0.1675(0.0047) 2.1029(0.5383)
50 0.0254(0.0047) 0.1375(0.0042) 2.1176(0.6382)
100 0.0209(0.0177) 0.0758(0.0200) 2.1351(0.5849)
150 0.0135(0.0491) 0.0402(0.0518) 2.0381(0.4749)
The average values and standard errorss of the root mean square geodesic errors for the models iWGPR, WGPR, and MGPR are presented in Tables 3 and 4 respectively for S3 and S4, offering important insights into the performance of each model under different scenarios. The experimental results clearly indicate that iWGPR achieves the lowest root mean square geodesic error. In contrast, consistent with findings on SPD manifolds, MGPR shows the highest root mean square geodesic error when the inherent geometric structure is neglected. Notably, even as the sampling points become sparser, iWGPR maintains relatively strong performance, whereas WGPR’s performance declines significantly.

### v2

@@ 2.3 Intrinsic Gaussian process regression
$\binom{y_{t^{*}}}{\boldsymbol{y}}\sim\mathcal{N}_{\mathcal{M}}\left(\binom{\mu_{t^{*}}}{\boldsymbol{\mu}},\left(\begin{array}[]{cc}\mathbf{E}^{\top}\boldsymbol{K}_{\mathbf{E},t^{*}}^{**}\mathbf{E},&\mathbf{E}^{\top}\boldsymbol{K}_{\mathbf{E},t^{*}}^{*}(\mathbf{E}\otimes\boldsymbol{I}_{n\times n})\\ (\mathbf{E}^{\top}\boldsymbol{K}_{\mathbf{E},t^{*}}^{*}(\mathbf{E}\otimes\boldsymbol{I}_{n\times n}))^{\top},&(\mathbf{E}\otimes\boldsymbol{I}_{n\times n})^{\top}(\boldsymbol{K}_{\mathbf{E},t^{*}}+\sigma^{2}\boldsymbol{I}_{nD\times nD})(\mathbf{E}\otimes\boldsymbol{I}_{n\times n})\end{array}\right)\right),$
where $\mathcal{N}_{\mathcal{M}}$ is the intrinsic Gaussian distribution. The remaining procedure for computing estimation and prediction is the same as the one in Subsection 2.3, where $(\mathbf{E}\otimes\boldsymbol{I}_{n\times n})^{\top}\boldsymbol{K}_{\mathbf{E},t^{*}}(\mathbf{E}\otimes\boldsymbol{I}_{n\times n})$ is replaced with $(\mathbf{E}\otimes\boldsymbol{I}_{n\times n})^{\top}(\boldsymbol{K}_{\mathbf{E},t^{*}}+\sigma^{2}\boldsymbol{I}_{nD\times nD})(\mathbf{E}\otimes\boldsymbol{I}_{n\times n})$. Theorem 2.8 below proves information consistency of the prediction obtained by the model (6). Suppose that the underlying true vector field $\boldsymbol{\gamma}(t)$ along the BPF $\mu(t)$ can be represented by real-valued function vectors $\boldsymbol{\gamma}_{\mathbf{E}}$ based on an orthogonal coordinate frame $\{\mathbf{E}_{t}\}_{t\in\mathbf{T}}$, where $\Gamma_{t\rightarrow s}(\mathbf{E}_{t})=\mathbf{E}_{s}$ for any $t,s\in\mathbf{T}$. Without loss of generality, we consider a special setting of $k_{\mathbf{E},t^{*}}$, i.e. the coordinates are independent of each other, i.e. $k_{\mathbf{E},t^{*}}\left(\cdot,\cdot\right)=\text{diag}(k_{\mathbb{R}^{Q}}\left(\cdot,\cdot,\boldsymbol{\theta}_{1}\right),\cdots,k_{\mathbb{R}^{Q}}\left(\cdot,\cdot,\boldsymbol{\theta}_{D}\right))$. Then, assume that the random element $\{\boldsymbol{\gamma}_{\mathbf{E},d}\},d=1,2,\cdots,D$ has a Gaussian process prior with mean zero and a bounded covariance function $k_{\mathbb{R}^{Q}}\left(\cdot,\cdot,\boldsymbol{\theta}_{d}\right)$ for any covariate $\boldsymbol{x}_{t}$. Suppose the $k_{\mathbb{R}^{Q}}\left(\cdot,\cdot,\boldsymbol{\theta}_{d}\right)$ is continuous in $\boldsymbol{\theta}_{d}$, further assume the estimator $\widehat{\boldsymbol{\theta}}_{d}$ converges to $\boldsymbol{\theta}_{d}$ as $n\rightarrow\infty$ almost surely, and the BPF $\mu(t)$ is known.
@@
Under the above assumptions, let the reproducing kernel Hilbert space norm be bounded and the expected regret term $E\left(\log\left|\boldsymbol{I}_{n\times n}+\sigma^{-2}\boldsymbol{K}_{d}\right|\right)=o(n)$ for $d=1,\cdots,D$, where $\boldsymbol{K}_{d}=\left(k_{\mathbb{R}^{Q}}\left(\boldsymbol{x}_{t_{i}},\boldsymbol{x}_{t_{j}};\boldsymbol{\theta}_{d}\right)\right)_{1\leq i,j\leq n}$ is the $d$-dimensional representation of the covariance matrix over $\{\boldsymbol{x}_{t_{i}}\}_{i=1}^{n}$ under the orthogonal coordinate frames $\{\mathbf{E}_{t}\}_{t\in\mathbf{T}}$, and $\sigma^{2}$ is the variance of the measurement error. Then, the estimator $\widehat{\boldsymbol{\gamma}}_{\mathbf{E},d}$ is information consistent to the true ${\boldsymbol{\gamma}}_{\mathbf{E},d}$, which means an estimator that guarantees convergence to the true parameter value as the sample size grows.
The proof is given in Section 6 of the Supplementary Materials. Furthermore, we also consider the consistency of the posterior distribution. This is shown in Theorem 2.9 and the proof is given in Section 7 of the Supplementary Materials.
@@
Suppose that the underlying true vector field $\boldsymbol{\gamma}(t)$ along the BPF $\mu(t)$ can be represented by real-valued function vectors $\boldsymbol{\gamma}_{\mathbf{E}}$ based on an orthogonal coordinate frame $\{\mathbf{E}_{t}\}_{t\in\mathbf{T}}$, where $\Gamma_{t\rightarrow s}(\mathbf{E}_{t})=\mathbf{E}_{s}$ for any $t,s\in\mathbf{T}$. Without loss of generality, we consider that the coordinates are independent of each other and that the BPF $\mu(t)$ is known. Let ${\boldsymbol{\gamma}}_{\mathbf{E},d}:\mathbb{R}^{Q}\rightarrow\mathbb{R}$ be a function and $\widehat{v}_{n,d}(\boldsymbol{x}_{t})$ denote the variance of the posterior distribution of the estimator $\widehat{\boldsymbol{\gamma}}_{E,d}$. Then, for every $d=1,2,\cdots,D$, we have $\sup_{\boldsymbol{x}\in\mathbb{R}^{Q}}\widehat{v}_{n,d}(\boldsymbol{x}_{t})\longrightarrow 0,\quad n\rightarrow\infty,$ almost surely monotonically, as well as in $L^{p}$ for every $p\in[1,\infty)$. Furthermore, if ${\boldsymbol{\gamma}}_{\mathbf{E},d}$ lies in the RKHS of $k_{\mathbb{R}^{Q}}\left(\cdot,\cdot,\boldsymbol{\theta}_{d}\right)$, for all $\boldsymbol{x}_{t}\in\mathbb{R}^{Q}$ and $d=1,2,\cdots,D$, we have $\widehat{\boldsymbol{\gamma}}_{\mathbf{E},d}(\boldsymbol{x}_{t})\xrightarrow{L^{2}}{\boldsymbol{\gamma}}_{\mathbf{E},d}(\boldsymbol{x}_{t}),\quad n\rightarrow\infty.$
@@
In Theorem 2.9, it needs the condition that $\mu(t)$ is known. But in practice, $\mu(t)$ is usually unknown and is estimated using data. In this case, we have the following conclusion. Consider the same conditions as in Theorem 2.9 and suppose $\sup_{t\in\mathcal{T}}d_{\mathcal{M}}^{2}(\widehat{\mu}(t),\mu(t))=O_{P}\left(n^{-1}\right)$ holds. Then, we have $(\Gamma_{\widehat{\mu}(t)\rightarrow\mu(t)}(\widehat{\boldsymbol{\gamma}}(\boldsymbol{x}_{t})))_{\mathbf{E},d}\xrightarrow{L^{2}}{\boldsymbol{\gamma}}_{\mathbf{E},d}(\boldsymbol{x}_{t})$, for all $\boldsymbol{x}_{t}\in\mathbb{R}^{Q}$, $n\rightarrow\infty$. Furthermore, from Theorem 6 of [23], assuming that the basepoint function is an intrinsic mean curve, condition $\sup_{t\in\mathcal{T}}d_{\mathcal{M}}^{2}(\widehat{\mu}(t),\mu(t))=O_{P}\left(n^{-1}\right)$ holds.
@@ 3 Simulation studies
In this section, we conduct several numerical experiments to illustrate the performance of the proposed iGPR model. We consider sample sizes of $N=30,50,100,150$ and two setups. The first is a random sampling scheme (Random), where 80% of the $N$ observation points are randomly selected as training data and the remaining are the test data. The second setup is a sorted sampling scheme (Sort), in which the last five samples of the $N$ sample points are designated as the test data and the remainings are the training data. All simulation results are based on 100 independent repetitions, and the average and standard errors of the root-mean-square geodesic errors are computed. To evaluate the performance of the proposed method, we compare it with wrapped Gaussian process regression (WGPR) [26] and vector-valued Gaussian process regression (MGPR) [1].
This section considers an example: the symmetric positive definite matrices space $\mathcal{S}_{++}^{3}$ with affine-invariant metric, while another example: the SPD manifold on the unit sphere, is presented in Section 9 of the Supplementary Materials. Let $\mu(t)$ be a geodesic curve $\mu(t)=\operatorname{Exp}(\mu(0),\operatorname{Log}(\mu(0),(2.4360,$ $-2.6465;-2.6465,8.4208)^{\top})t),$ where $\mu(0)=(1,0;0,1)$, the time domain is $t\in[0,1]$ and the predictor is time, i.e, $\boldsymbol{x}_{t}=t$. The initial frame $(1,0;0,0)$, $(0,1;1,0)$ and $(0,0;0,1)$ are utilized in the tangent space of $\mu(0)$. Let the coordinates be independent using the tangle space of $\mu(0)$ as initial tangle space to generate the covariance function. We consider the following two scenarios. (S1) We use the following Gaussian kernel function matrix as the intrinsic covariance function under the frame $\mathbf{E}$,
$k_{\mathbf{E},t^{*}}(\boldsymbol{x},\widetilde{\boldsymbol{x}}):=\begin{pmatrix}\exp({\frac{-\|\boldsymbol{x}-\widetilde{\boldsymbol{x}}\|_{2}^{2}}{\theta_{j1}}})&0&0\\ 0&\exp({\frac{-\|\boldsymbol{x}-\widetilde{\boldsymbol{x}}\|_{2}^{2}}{\theta_{j2}}})&0\\ 0&0&\exp({\frac{-\|\boldsymbol{x}-\widetilde{\boldsymbol{x}}\|_{2}^{2}}{\theta_{j3}}})\end{pmatrix},\boldsymbol{x},\widetilde{\boldsymbol{x}}\in[0,1];j=1,2,3,$
Table 1: The average and standard errors of the root mean square geodesic errors of WGPR, iGPR and MGPR on the SPD manifold $\mathcal{S}_{++}^{3}$ with S1.
Scheme $\boldsymbol{\theta}$ $n$ iGPR WGPR MGPR
@@
Average values and standard errors of the root mean square geodesic errors from three methods, iGPR, WGPR and MGPR, are shown in Tables 1-2, corresponding to scenarios S1 and S2 respectively. From both Tables, we see that iGPR has the smallest root mean square geodesic errors, indicating that iGPR has a significant improvement in predictive performance compared to WGPR and MGPR. In addition, MGPR has the largest root mean square geodesic errors, as it fails to account for the geometric information inherent in manifold-valued data during modeling, thus violating the intrinsic structure of the manifold. Although WGPR incorporates the local nonlinear structure of the manifold through exponential and logarithmic maps, it still disrupts the manifold’s intrinsic geometric structure by directly computing tangent vectors from different tangent spaces during modeling. As the sampling points become sparser, the proposed method still maintains excellent performance, whereas the performance of WGPR declines significantly.
We also investigate the ability of the proposed method to reconstruct the covariance function of Gaussian process. The results are shown in Section 8 of the Supplementary Materials, which illustrates the estimated covariance function in the tangent space compared to the true covariance function with kernel parameters for covariance structure reflecting the mean values from 100 experiments. The heatmaps indicate that the estimated covariance functions align more closely with the true one compared with WGPR. Additionally, similar experiments were conducted on the unit sphere, and the results are generally consistent with those obtained on the SPD manifold. For more details, see Section 9 of the Supplementary Materials.
Table 2: The average and standard errors of the root mean square geodesic errors of WGPR, iGPR and MGPR on the SPD manifold $\mathcal{S}_{++}^{3}$ with S2.
Scheme $\boldsymbol{\sigma^{2}}$ $n$ iGPR WGPR MGPR